OpenStudy (anonymous):

One earthquake has magnitude 4.5. If a second earthquake has 760 times as much energy as the first, find the magnitude of the second quake.

OpenStudy (anonymous):

S = 760*10^6.75 M = 2/3log (760 * 10^6.75) M = 2/3log(760*10^6.75)

OpenStudy (anonymous):

I need help to solve this equation to get the magnitude of the second quake

OpenStudy (anonymous):

The math book has an example of M = 2/3log(2*10^4.5) is approx 3.201 but how did they get that answer?

OpenStudy (anonymous):

If I can figure that out, I should be able to solve for M

ganeshie8 (ganeshie8):

how did u get 6.75 ?

OpenStudy (anonymous):

4.5 = 2/3log(S/So) 4.5(3/2) = log(S/So) 6.75 = log(S/So)

OpenStudy (anonymous):

Is there some formula to calculate magnitude of any earthquake

OpenStudy (anonymous):

Difference in orders of magnitude=log(a/b) Moment magnitude scale=2/3log(s/So)

OpenStudy (anonymous):

I used the moment magnitude scale for this problem